# Consensus-based construction of high-dimensional free energy surface

### v3

@@ Conclusion
We have presented a consensus-based approach to construct high-dimensional FESs by reformulating the construction task as a minimax optimization problem. Rather than seeking the direct fitting, this method essentially establishes an adversarial learning of FESs by simultaneously optimizing the target function approximation and the training set. While the common approaches mainly focus on the efficient exploration of the phase space in the presence of local minima, the present method further accounts for the discretization error that has been broadly overlooked in FES construction. Adaptive sampling of the max-residue regime is achieved through the consensus-based sampling of a posteriori residue-induced distribution in the form of a stochastic particle system in the CV space. Given the fact the sampling only relies on the first and second-moment estimation, the method could be particularly efficient for high-dimensional problems. While the numerical results of biomolecular systems have demonstrated the effectiveness for FES construction, the present framework of unifying the residual minimization and max-residue enhanced sampling is quite general for model reduction of complex systems, e.g., the stochastic reduced dynamics (e.g., see Ref. [61]) with a state-dependent memory for multiscale physical systems.
Acknowledgements
The work is supported in part by the National Science Foundation under Grant DMS-2143739 and the ACCESS program through allocation MTH210005.
Appendix A Proof of Proposition 2
The sampling distribution is governed by the following McKean stochastic differential equation
$\dot{\mathbf{z}}^{i}_{t}=-\frac{1}{\gamma}\nabla_{\mathbf{z}}G(\mathbf{z}^{i}_{t})+\sqrt{\frac{2}{\kappa_{h}\gamma}}\mathbf{\xi}_{i}(t),$ (16)
where $G(\mathbf{z}_{t})=\frac{1}{2}(\mathbf{z}_{t}-\mathbf{m}_{t})^{T}V_{t}^{-1}(\mathbf{z}_{t}-\mathbf{m}_{t})$ and $\gamma$ is the friction coefficient. We want to show it converges to a steady state as an approximation of $q^{\ast}$ with the proper choices of $\mathbf{m}_{t}$ and $V_{t}$ given by
$\mathbf{m}_{t}$ $=\sum_{i=1}^{N_{w}}\mathbf{z}^{i}_{t}\hat{p}(\mathbf{z}^{i}_{t}),$ (17)
$V_{t}$ $=\kappa_{t}\sum_{i=1}^{N_{w}}(\mathbf{z}^{i}_{t}-\mathbf{m}_{t})(\mathbf{z}^{i}_{t}-\mathbf{m}_{t})^{T}\hat{p}(\mathbf{z}^{i}_{t}),$
where $\hat{p}(\mathbf{z})=\frac{\exp{(-\kappa_{l}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))}}{\sum_{i=1}^{N_{w}}\exp{(-\kappa_{l}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}^{i})})}$ and $\kappa_{t}=\kappa_{l}+\kappa_{h}$.
Proposition 2. Suppose $\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z})$ takes a local quadratic approximation in form of $\frac{1}{2}(\mathbf{z}-\boldsymbol{\mu})^{T}\Sigma^{-1}(\mathbf{z}-\boldsymbol{\mu})$, $q_{t}\to\frac{\exp{(-\kappa_{h}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))}}{\int\exp{(-\kappa_{h}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z}))}\mathrm{d}\mathbf{z}}$ as $t\to\infty$, given $\kappa_{t}=\kappa_{l}+\kappa_{h}$.
Proof.
Let $q_{\infty}(\mathbf{z})$ denote the invariant distribution of the McKean SDE (16). Then $q_{\infty}(\mathbf{z})$ must be the invariant distribution of the following McKean stochastic dynamics
$\dot{\mathbf{z}}=-\frac{1}{\gamma}V^{-1}_{\kappa_{l},\infty}(\mathbf{z}-\mathbf{m}_{\kappa_{l},\infty})+\sqrt{\frac{2}{\gamma\kappa_{h}}}\xi(t),$ (18)
Since $\mathcal{L}_{\mathcal{N}}(\mathbf{z})=\frac{1}{2}(\mathbf{z}-\boldsymbol{\mu})^{T}\Sigma^{-1}(\mathbf{z}-\boldsymbol{\mu})$ is quadratic, the re-weighted density of a Gaussian distribution $q(\mathbf{z})\sim\mathcal{N}(\mathbf{m},V)$ remains Gaussian, i.e., $q(\mathbf{z})e^{-\kappa_{l}\mathcal{L}_{\mathcal{N}}^{-}(\mathbf{z})}\propto~{}\mathcal{N}({\mathbf{m}}_{\kappa_{l}},{V}_{\kappa_{l}})$, where ${\mathbf{m}}_{\kappa_{l}}$ and ${V}_{\kappa_{l}}$ are defined by
$\begin{split}\mathbf{m}_{\kappa_{l}}&=(V^{-1}+\kappa_{l}\Sigma^{-1})^{-1}(\kappa_{l}\Sigma^{-1}\boldsymbol{\mu}+V^{-1}\mathbf{m}),\\ V_{\kappa_{l}}&=(V^{-1}+\kappa_{l}\Sigma^{-1})^{-1}.\end{split}$ (19)
In particular, we choose $\mathbf{m}=\mathbf{m}_{\kappa_{l},\infty}$ and $V=\kappa_{h}^{-1}V_{\kappa_{l},\infty}$, then Eq. (19) yields
$\begin{split}\mathbf{m}_{\kappa_{l},\infty}&=(\kappa_{h}V_{\kappa_{l},\infty}^{-1}+\kappa_{l}\Sigma^{-1})^{-1}(\kappa_{l}\Sigma^{-1}\boldsymbol{\mu}+\kappa_{h}V_{\kappa_{l},\infty}^{-1}\mathbf{m}_{\kappa_{l},\infty}),\\ \kappa_{t}^{-1}V_{\kappa_{l},\infty}&=(\kappa_{h}V_{\kappa_{l},\infty}^{-1}+\kappa_{l}\Sigma^{-1})^{-1}.\end{split}$
It is easy to show that by choosing $\kappa_{t}=\kappa_{l}+\kappa_{h}$, $\mathbf{m}_{\kappa_{l},\infty}$ and $V_{\kappa_{l},\infty}$ recovers $\boldsymbol{\mu}$ and $\Sigma$, respectively, and the invariant density takes the form
$q_{\infty}(\mathbf{z})\sim\mathcal{N}\left(\boldsymbol{\mu},\kappa_{h}^{-1}\Sigma\right).$
Appendix B Simulation setup
@@
All the MD simulations are performed using the package GROMACS 2019.2 [62]. The simulation is carried out on Intel(R) Xeon(R) Platinum 8260 CPU.
B.1 ala2 (the two-dimensional problem)
@@
The Ace-Ala-Nme (ala2) molecule is modeled by the Amber99SB force field [58]. The molecules are dissolved in 383 water molecules in a periodic simulation cell. The cut-off radius of the van der Waals interaction is 0.9 nm. The Coulomb interaction is treated with the smooth particle mesh Ewald method with a real space cutoff of 0.9 nm and a reciprocal space grid spacing of 0.12 nm. The system is integrated with the leap-frog scheme at time step 2 fs. The temperature of the system is set to 300 K by a velocity-rescale thermostat [63] with a relaxation time of 0.2 ps. The Parrinello-Rahman barostat [64] with a relaxation time scale of 1.5 ps and a compressibility of $4.5\times 10^{-5}$ bar-1 is coupled to the system to control the pressure to 1 bar. The hydrogen atom is constrained by the LINCS algorithm[65] and the H–O bond and H–O–H angle of water molecules are constrained by the SETTLE algorithm[66].
B.2 s1pe (three-dimensional problem)
The s-(1)-phenylethyl (s1pe) molecule is modeled by the CHARMM general force field (CGenFF) [59]. The molecule is dissolved in 546 TIP3P water molecules in a (2.69 nm)3 dodecahedron box. The cut-off radius of the van der Waals interaction is 1 nm. The Coulomb interaction is treated with the smooth particle mesh Ewald method with a real space cutoff of 1 nm and a reciprocal space grid spacing of 0.12 nm. The system is integrated with the leap-frog scheme at time step 2 fs. The temperature of the system is set to 298 K by a velocity-rescale thermostat [63] with a relaxation time of 0.2 ps. The Parrinello-Rahman barostat pressure couple, the hydrogen atom constraint, the H–O bond constraint, and the H–O–H angle of water molecules constraint are the same as the previous one.
B.3 (s1pe)3 (the nine-dimensional problem)
@@
The (s1pe)3 molecule is modeled by the CHARMM general force field (CGenFF) [59]. The molecule is solvated in a (4.2 nm)3 dodecahedron box with 1622 TIP3P water molecules. Other simulation setups are similar to the s1pe molecule.
B.4 ala16 (the thirty-dimensional problem)
@@
The Ace-(Ala)15-Nme (ala16) molecule is modeled by the Amber99SB force field [58]. The molecule is solvated in 2258 water molecules in a (4.62 nm)3 dodecahedron box. Other simulation setups are similar to the ala2 molecule.
Appendix C Traning
@@
The training data is collected during the sampling process by restricted dynamics. The initial $10\%$ steps of the restricted dynamics are used as equilibrium and the rest $90\%$ steps are used to calculate the mean force. The FES are parameterized as a fully connected neural network. The depth and width of the NN are shown in Table 3. The NNs are trained by Adam for $100000$ steps with a learning rate $1\times 10^{-3}$. For each training step, $5000$ sampling points are randomly selected from the data set. All the training process is carried out on v100 with 32768 MB memory.
@@
Appendix D Additional result
D.1 Additional result of the three-dimensional FES (s1pe)
Figure 7-9 shows the additional 2D projections of the 3D FES (molecule s1pe) constructed by the Rid and presented CES method. For each case, the reference is constructed as a 2D FES using the metadynamics. The CES method yields a better agreement with the reference.
Figure 7: The 3D FES for the s1pe molecule projected on the $\phi-\psi$ plane with the third variable $\omega=-1$ (first row) and $\omega=-2$ (second row). (a) 2D FES by metadynamics (reference) (b) RiD (c) CES.
Figure 8: The 3D FES of the s1pe molecule on the $\omega-\psi$ plane with the third variable $\phi=0$ (first row) and $\phi=-1$ (second row). (a) 2D FES by metadynamics (reference) (b) RiD (c) CES.
Figure 9: The 3D FES of the s1pe molecule on the $\omega-\phi$ plane with the third variable $\psi=0$ (first row) and $\psi=-1$ (second row). (a) 2D FES by metadynamics (reference) (b) RiD (c) CES.
D.2 Additional results of the nine-dimensional FES (s1pe)3
The addition 2D projections of the 9D FES for molecule (s1pe)3 are presented in Figure 10. Similar to the 3D case, the FES constructed by the present CES method shows a better agreement with the reference constructed as a 2D FES using metadynamics.
Figure 10: The 9D FES for the (s1pe)3 molecule projected on the $\omega_{1}-\phi_{1},\omega_{2}-\phi_{2},\omega_{2}-\psi_{2},\phi_{2}-\psi_{2},\omega_{3}-\phi_{3},\omega_{3}-\psi_{3},\phi_{3}-\psi_{3}$ plane from top to bottom. (a) 2D FES by metadynamics (reference) (b) RiD (c) CES.
D.3 Additional results of the thirty-dimensional FES (ala16)
The addition 2D projections of the 30D FES for molecule ala16 are presented in Figure 11. We note that the projection on the $\phi_{5}-\psi_{5}$ plane is significantly different from other projections such as the $\phi_{1}-\psi_{1}$ plane in the main context. While the ala16 molecule consists of 15 sequential alanine residues, the FES for individual $\phi-\psi$ projections shows different features. The numerical results of the present CES method show good agreement with the reference.
Figure 11: The 30D FES of the ala16 projected on the $\phi_{5}-\psi_{5}$ plane. (a) 2D FES constructed by metadynamics (reference) (b) 30D FES constructed by the CES method.
Appendix E Code availablity

### v4

@@ 4 Conclusions
We developed a consensus-based sampling approach for adaptive surrogate construction in systems with complex energy landscapes. The method reformulates the construction task as a minimax optimization problem by simultaneously optimizing the target function approximation and the training set. A unique feature of the method is its ability to unify the dynamical exploration of high-dimensional phase space in the presence of prevalent energy barriers with efficient sampling adaptive to the posterior residual error. As a motivating example, we considered the long-standing problem of constructing high-dimensional FESs for MD systems. Compared with existing approaches [41] focusing on various enhanced sampling strategies to overcome energy barriers, the present method further enables the residual-based adaptive sampling that remains an open problem in FES construction. On the other hand, unlike the common adaptive sampling strategies based on the generative model [43] and important sampling [16] used for solving high-dimensional PDEs, the present method achieves efficient sampling of the max-residual region in the form of an interacting particle system that does not rely on the free query of the arbitrary point in the phase space and the gradient computation of the residual function. Given the fact the sampling only relies on the first and second-moment estimation, the method could be particularly efficient for high-dimensional problems. While the numerical results of biomolecular systems have demonstrated the effectiveness for the high-dimensional FES construction, the present method provides a framework for unifying the phase space exploration and residual-based adaptive sampling, and can be broadly applied to multiscale systems involving constrained sampling dynamics and surrogate modeling, such as rare event sampling [13], stochastic reduced dynamics [31, 17] and uncertainty quantification in high-dimensional constrained phase space [29].
Appendix A Simulation setup
@@
All the MD simulations are performed using the package GROMACS 2019.2 [30] and open-source, community-developed PLUMED library [45]. The simulation is carried out on Intel(R) Xeon(R) Platinum 8260 CPU.
A.1 ala2
@@
The Ace-Ala-Nme (ala2) molecule is modeled by the Amber99SB force field [22]. The molecule is solvated in an aqueous environment with 383 TIP3P water molecules. Periodic boundary conditions are imposed along each direction. The cut-off radius of the van der Waals interaction is 0.9 nm. The long-range Coulomb interaction is treated with the smooth particle mesh Ewald method with a real space cutoff of 0.9 nm and a reciprocal space grid spacing of 0.12 nm. The system is integrated with the leap-frog scheme at time step 2 fs. The temperature is set to be 300 K by a velocity-rescale thermostat [5] with a relaxation time of 0.2 ps. The Parrinello-Rahman barostat [36] with a relaxation time scale of 1.5 ps and a compressibility of $4.5\times 10^{-5}$ bar-1 is coupled to the system to control the pressure to 1 bar. The hydrogen atom is constrained by the LINCS algorithm [21] and the H–O bond and H–O–H angle of water molecules are constrained by the SETTLE algorithm [33].
A.2 s1pe
@@
The s-(1)-phenylethyl (s1pe) molecule is modeled by the CHARMM general force field (CGenFF) [50]. The molecule is dissolved in 546 TIP3P water molecules in a (2.69 nm)3 dodecahedron box. The cut-off radius of the van der Waals interaction is 1 nm. Other setups are similar to the ala2 molecule.
A.3 (s1pe)3
@@
The (s1pe)3 molecule is modeled by the CHARMM general force field (CGenFF) [50]. The molecule is solvated in a (4.2 nm)3 dodecahedron box with 1622 TIP3P water molecules. Other setups are similar to the ala2 molecule.
A.4 ala16
The Ace-(Ala)15-Nme (ala16) molecule is modeled by the Amber99SB force field [22]. The molecule is solvated in a (4.62 nm)3 dodecahedron box with 2258 TIP3P water molecules. Other setups are similar to the ala2 molecule.
Appendix B Training
@@
The training data is collected by restrained molecular dynamics. The initial $10\%$ steps of the restrained dynamics are used as equilibrium and the rest $90\%$ steps are used to calculate the mean force. The FES are parameterized as a fully connected neural network. The depth and width of the NN are shown in Table 2. The NNs are trained by Adam for $100000$ steps with a learning rate $1\times 10^{-3}$. For each training step, $5000$ sampling points are randomly selected from the data set. All the training processes are conducted using an Nvidia GPU V100 with 32GB memory.
@@
Appendix C Additional numerical results
C.1 1D Rastrigin function
[…]
Table 3: The max-residual point $z^{*}_{i}$ and its second derivative $f^{\prime\prime}(z^{*}_{i})$, as well as their numerical estimations from the present sampling method at each iteration. Specifically, the max-residual point $z^{*}_{i}$ can be estimated by $m_{i}$ and the second derivative $f^{\prime\prime}(z^{*}_{i})$ can be estimated by $f^{\prime\prime}_{\theta_{i}}\pm{V}^{-1}$.
$i$ 0 1 2 3 4 5 6
$z^{*}_{i}$ 0.000 2.000 -2.000 0.996 -0.996 2.805 -2.806
$m_{i}$0.000 2.000 -2.000 0.996 -0.995 2.805 -2.806
$f^{\prime\prime}(z^{*}_{i})$ 41.478 41.478 41.478 41.468 41.462 15.273 15.539
$f^{\prime\prime}_{\theta_{i}}\pm{V}^{-1}$ 40.974 37.442 36.916 38.283 39.519 11.841 13.380
$i$ 6 7 8 9 10 11 12
$z^{*}_{i}$2.806 1.500 -1.502 0.503 -0.501 2.416 -2.413
$m_{i}$2.806 1.499 -1.503 0.502 -0.501 2.414 -2.413
$f^{\prime\prime}(z^{*}_{i})$ 15.539 -37.478 -37.473 -37.474 -37.478 -31.901 -31.688
$f^{\prime\prime}_{\theta_{i}}\pm{V}_{i}^{-1}$ 15.380 -36.796 -36.271 -37.515 -37.285 -31.521 -31.189
C.2 The 2D FES for molecule ala2
Tab.  4 shows the approximation error of the constructed FES surrogate at each iteration. The total computational cost is presented in Tab. 1 in the main manuscript.
Table 4: The $l_{2}$ and $l_{\infty}$ error of the FES surrogate by the present CAS method for each sampling iteration.
iteration 1 2 3 4 5 6 7
$l_{\infty}$ error 35.21 30.62 22.38 16.41 14.83 9.86 10.68
$l_{2}$ error 9.82 7.52 6.84 3.77 2.36 3.37 1.88
C.3 The 3D FES for molecule s1pe
Fig. 7-9 shows the additional 2D projections of the 3D FES (molecule s1pe) constructed by the Rid and presented CAS method. For each case, the reference is constructed as a 2D FES using the metadynamics. The computational cost and accuracy of CAS and RiD are shown in Tab. 5. It shows that the CAS method achieves better accuracy with less computational cost.
Table 5: The accuracy of the constructed 3D FES (the s1pe molecule) and computational time for RiD and the present CAS method. The $l_{2}$ and $l_{\infty}$ error are calculated up to $40$ KJ/mol. For each case, the FES is projected onto a 2D plane with the third variable fixed; the reference solution is constructed as a 2D FES by the metadynamics. The simulation time of the CAS method is multiplied by $20$ since 20 walkers are used.
Method Restraint Accuracy Time
$l_{2}$ error $l_{\infty}$ error Sampling Train
RiD $\psi=1.5$ $5.76$ $25.72$ $423.33$ $8$ (GPU)
$\omega=1.5$ $12.04$ $49.13$
CAS $\psi=1.5$ $2.44$ $11.21$ $4.81\times 20$ $0.84$ (GPU)
$\omega=1.5$ $3.89$ $28.80$
C.4 The 9D FES for molecule (s1pe)3
The addition 2D projections of the 9D FES for molecule (s1pe)3 are presented in Fig. 10. Similar to the 3D case, the FES constructed by the present CAS method shows a better agreement with the reference constructed as a 2D FES using metadynamics.
C.5 The 30D FES for molecule ala16
The addition 2D projections of the 30D FES for molecule ala16 are presented in Fig. 11. We note that the projection on the $\phi_{5}-\psi_{5}$ plane is significantly different from other projections such as the $\phi_{1}-\psi_{1}$ plane in the main context. While the ala16 molecule consists of 15 sequential alanine residues, the FES for individual $\phi-\psi$ projections shows different features. The numerical results of the present CAS method show good agreement with the reference.
Appendix D Code availablity
@@
Figure 7: The 3D FES for the s1pe molecule projected on the $\phi-\psi$ plane with the third variable $\omega=-1$ (first row) and $\omega=-2$ (second row). (a) 2D FES by metadynamics (reference) (b) RiD (c) CAS.
Figure 8: The 3D FES of the s1pe molecule on the $\omega-\psi$ plane with the third variable $\phi=0$ (first row) and $\phi=-1$ (second row). (a) 2D FES by metadynamics (reference) (b) RiD (c) CAS.
Figure 9: The 3D FES of the s1pe molecule on the $\omega-\phi$ plane with the third variable $\psi=0$ (first row) and $\psi=-1$ (second row). (a) 2D FES by metadynamics (reference) (b) RiD (c) CAS.
Figure 10: The 9D FES for the (s1pe)3 molecule projected on the $\omega_{1}-\phi_{1},\omega_{2}-\phi_{2},\omega_{2}-\psi_{2},\phi_{2}-\psi_{2},\omega_{3}-\phi_{3},\omega_{3}-\psi_{3},\phi_{3}-\psi_{3}$ plane from top to bottom. (a) 2D FES by metadynamics (reference) (b) RiD (c) CAS.
Figure 11: The 30D FES of the ala16 projected on the $\phi_{5}-\psi_{5}$ plane. (a) 2D FES constructed by metadynamics (reference) (b) 30D FES constructed by the present CAS method.
Acknowledgments
The work is supported in part by the National Science Foundation under Grant DMS-2110981 and the ACCESS program through allocation MTH210005.